pH of 10 power -9M HCl is approx.? ( not getting correct ans.)

Grade:12

2 Answers

Naveen Kumar
9 years ago
Here H+ obtained by only HCl is =10-9M
As we also know that the H+ obtained by dissociation of water=10-7M
So here we can see thaty the H+ released by water is more that the dissociation of HCl
So consider both the H+ and caculate pH
Total H+=10-7.+10-9 M
pH=-logH+
In general, If the concentration of acid is quite low and the H+ released by the aciid is compareble to 10-7M(obtained by water),i.e(consider upto 100times, for example if acid is giving 10-5M or below it) then consider both H+ while calculatingf pH)
Hope you would get it.
Shriya Mehrotra
9 years ago
thanku sir.